OpenStudy (anonymous):

What is the value of (8 – 5)!? A. 5040 B. 24 C. 6 D. 40,200

OpenStudy (anonymous):

man i had it...

OpenStudy (mathstudent55):

Do you know what x! means?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

You know about factorials

OpenStudy (anonymous):

no

OpenStudy (mathstudent55):

The exclamation mark is the symbol for "factorial". Do you know what factorial means?

OpenStudy (anonymous):

no I dont

OpenStudy (anonymous):

n! = n(n-1)(n-2)......1, For example, 4! = 4 x 3 x 2 x 1 = 24

OpenStudy (anonymous):

6

OpenStudy (anonymous):

Yep..!!

OpenStudy (anonymous):

thank you

OpenStudy (mathstudent55):

It's a simple concept. A factorial number is a number multiplied by one less than itself, then multiplied by one less than that until you multiply it by 1. It's easier to understand with an example. 5! = 5 * 4 * 3 * 2 * 1 = 120 3! = 3 * 2 * 1 = 6

OpenStudy (anonymous):

How many combinations of 3 students can a teacher choose from 32 students? A. 96 B. 4960 C. 29,760 D. 32,768

OpenStudy (mathstudent55):

In your case, first do the subtraction inside the parentheses. Then do the result of the subtraction factorial. For example: (12 - 7)! = 5! = 5 * 4 * 3 * 2 * 1 = 120

OpenStudy (anonymous):

From 32 people, you have to choose 3 people which can be done in 32C3 ways. Just calculate the answer

OpenStudy (mathstudent55):

\(_nC_r = \dfrac{n!}{r!(n - r)!}\) \(_{32}C_3 \) \(= \dfrac{32!}{3!(32 - 3)!}\) \( = \dfrac{32 \cdot 31 \cdot 30 \cdot 29!}{3!29!}\) \(= \dfrac{32 \cdot 31 \cdot 30 \cdot \cancel{29!}}{3 \cdot 2 \cdot\cancel{29!}}\) \(= 32 \cdot 31 \cdot 5 \) \(= 4960\)